OpenStudy (anonymous):

What is the relationship between the diameter of a circle and the area of the circle? My answer: pi x (0.5d)^2 Is this correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

2pir and pir^2

OpenStudy (anonymous):

diameter = 2*r, area=pi*r^2.area/diameter=pi*r^2/2r=pi*r/2